Ask
your own question, for FREE!
Probability
10 Online
Sampling without replacement: An urn contains 2 black and 3 white balls. 2 balls are drawn at random without replacement. The probability that the sample contains at most one black ball. Supposedly the answer is .90, but I am not sure how to get that.
Still Need Help?
Join the QuestionCove community and study together with friends!
oooh AT MOST
compute the probability that there are no black balls compute the probability that there is one black ball then add
Probability No Black \[ \frac 3 5 \frac 2 4=\frac 3 {10} \] Probability One Black \[ \frac{2\ 3}{5\ 4}+\frac{3\ 2}{5\ 4}=\frac 3 5 \] Add them you obtain .9
@satellite73 the tree that I made is |dw:1397620597682:dw| P( no black): P(WW)= (3/5)(2/4) --> .3 P( 1 back): P(BW) + P(WB)= (2/5)(2/4) + (3/5)(2/4) --> .5 .3 + .5= 0.80 not 0.90
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
9 hours ago
1 Reply
0 Medals
5 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal